Accelerator Update for April 22, 2009 - April 24, 2009 |
Wednesday April 22 The day shift began with Operations monitoring the Tevatron (TeV) store 6991 (160.22E30) and the Recycler stash (160.1E10), with the Antiproton Source stacking (11.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. NuMI and MiniBooNE users took protons throughout the day shift. MTest users took beam most of the day, but not from 2:15 PM to 3:45 PM due to LCW cooling problems for beamline power supplies. Operations disabled the MTest experiment at 6 PM. Operations terminated store 6991 at 6:30 PM and put the TeV into shot setup. Operations established store 6992 at 8 PM with an initial luminosity of 337.68E30. Antiprotons were stacked and stashed in the Recycler until 10:45 PM; the Recycler injectionn kicker coolant pump failed. NuMI and MiniBooNE users took beam until 11 PM when Operations halted beam to prepare Main Injector for an access. Technicians accessed the Main Injector to repair the coolant pump at 11:30 PM. Thursday April 23 The midnight shift began with Operations monitoring store 6992 (177.75E30) and the Pbar stack (51.4E10), with MI in access to repair a Recycler kicker pump. The Main Injector access for the Recycler kicker repair lasted until 4 AM. Operations restored beam to Pbar, NuMI, and MiniBooNE by 4:45 AM. Operations established beam to MTest users at 7:30 AM. NuMI and MiniBooNE users took protons during the day shift with only minor interruptions. MTest users took beam until 3 PM when they accessed their hall. Operations terminated store 6992 at 4:30 PM. Operations established store 6993 at 6:30 PM with an initial luminosity of 288.46E30. Antiprotons were stacked and stashed in Recycler all evening except from 8:45 PM to 9:30 PM. Operators investigated a false alarm indication. NuMI and MiniBooNE users took protons throughout the evening shift except during the alarm investigation. MiniBooNE also saw limited beam from 10:30 PM to 11:00 PM due to beam position problems. Friday April 24 The midnight shift began with Operations monitoring store 6993 (128.43E30) and the stash (162.6E10), with Pbar stacking (19.4E10), and with NuMI and MiniBooNE taking beam. Antiprotons were stacked and stashed in Recycler and NuMI and MiniBooNE took protons throughout the night shift. Operations established beam to MTest at 7 AM. The Plans for Friday and the Weekend The plans for today are to terminate the store and allow machine accesses. The plans for the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/27/2009 email Fermilab |
| Security, Privacy, Legal | |